OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6 Answer slope = -5; y-intercept = (0, -6) slope = ; y-intercept = (0, 2) slope = - ; y-intercept = (0, 2) slope = ; y-intercept = (0, -6)

OpenStudy (anonymous):

a, b, c, or d

OpenStudy (precal):

5x-3y=-6 -3y=-5x-6 y=(-5x-6)/-3 y=(5/3)x+2 slope is (5/3) y-intercept is 2